Solve the equation. y′′−6y′+9y=0 ;  y(0)=2,y′(0)=−4

a. using differential equation method

b. using Laplace transformation method


1
Expert's answer
2021-07-29T16:46:27-0400

a. Characteristic equation:

"{k^2} - 6k + 9 = 0"

"{(k - 3)^2} = 0"

"{k_1} = {k_2} = 3"

Then

"y = {C_1}{e^{3x}} + x{C_2}{e^{3x}} \Rightarrow y' = 3{C_1}{e^{3x}} + {C_2}{e^{3x}} + 3x{C_2}{e^{3x}}"

"y(0) = 2,\,y'(0) = - 4 \Rightarrow \left\{ {\begin{matrix}
{{C_1} = 2}\\
{3{C_1} + {C_2} = - 4}
\end{matrix}} \right. \Rightarrow {C_1} = 2,\,{C_2} = - 10"

Then

"y = 2{e^{3x}} - 10x{e^{3x}}"

Answer: "y = 2{e^{3x}} - 10x{e^{3x}}"

b. Moving from originals to images:

"y \to Y"

"y' \to pY - y(0) = pY - 2"

"y'' \to {p^2}Y - py(0) - y'(0) = {p^2}Y - 2p + 4"

Then

"{p^2}Y - 2p + 4 - 6\left( {pY - 2} \right) + 9Y = 0"

"Y\left( {{p^2} - 6p + 9} \right) = 2p - 4 - 12"

"Y = \frac{{2p - 16}}{{{p^2} - 6p + 9}} = \frac{{2p - 16}}{{{{(p - 3)}^2}}} = \frac{2}{{p - 3}} - \frac{{10}}{{{{(p - 3)}^2}}}"

"Y \leftarrow y"

"\frac{1}{{p - 3}} \leftarrow {e^{3x}}"

"\frac{1}{{{{\left( {p - 3} \right)}^2}}} \leftarrow x{e^{3x}}"

Then

"y = 2{e^{3x}} - 10x{e^{3x}}"

Answer: "y = 2{e^{3x}} - 10x{e^{3x}}"